# Consensus-based construction of high-dimensional free energy surface

### v3

@@
$\lim_{\kappa\to\infty}\left(-\frac{1}{\kappa}\log\left(\int\exp{(-\kappa f(\mathbf{z}))}\mathrm{d}\rho^{*}(\mathbf{z})\right)\right)=f(\mathbf{z}^{*})$ (9)
@@
holds true for any compactly supported probability measure $\rho^{\ast}$, where $\mathbf{z}^{\ast}\in\operatorname{supp}(\rho^{*})$ uniquely minimizes the function $f$. This enables us to identify the max-residual point from a collection of samples $\left\{\mathbf{z}^{i}\right\}_{i=1}^{N_{w}}$ by the first-order momentum $\mathbf{m}$ under the weighted density function $p(\mathbf{z})$, i.e.,
@@
$\mathbf{m}=\int\mathbf{z}p(\mathbf{z})\mathrm{d}\mathbf{z}\approx\sum_{i=1}^{N_{w}}\mathbf{z}^{i}\hat{p}(\mathbf{z}^{i})\quad\hat{p}(\mathbf{z})=\frac{\exp{(-\kappa_{l}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))}}{\sum_{i=1}^{N_{w}}\exp{(-\kappa_{l}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}^{i})})},$ (10)
where $\kappa_{l}^{-1}$ represents a low temperature limit. However, the integration is subject to the so-called curse of dimensionality as the number of CVs increases. Instead, we treat sampler $\mathbf{z}^{i}$ as a random walker $\mathbf{z}^{i}_{t}$ governed by the following McKean stochastic differential equation
@@
$\dot{\mathbf{z}}^{i}_{t}=-\frac{1}{\gamma}\nabla_{\mathbf{z}}G(\mathbf{z}^{i}_{t})+\sqrt{\frac{2}{\kappa_{h}\gamma}}\mathbf{\xi}_{i}(t),$ (11)
where $G(\mathbf{z}_{t})=\frac{1}{2}(\mathbf{z}_{t}-\mathbf{m}_{t})^{T}V_{t}^{-1}(\mathbf{z}_{t}-\mathbf{m}_{t})$ denotes an adaptively constructed conservative potential function and $\gamma$ is the friction coefficient. The formulations of $\mathbf{m}_{t},V_{t}$ are specified in (12) with the rationale discussed in the next section. Consequently, $G(\mathbf{z}_{t})$ navigates the random walkers (i.e., individual particles) towards $\mathbf{m}_{t}$, which represents the region of large residual error. The second term in Eq. (11) represents a stochastic term where $\gamma$ represents the friction coefficient and $\mathbf{\xi}(t)$ represents the standard Gaussian white noise characterized by zero mean and covariance $\mathbb{E}[\mathbf{\xi}_{i}(t)\mathbf{\xi}_{j}(t^{\prime})]=\delta_{ij}\delta(t-t^{\prime})$.
@@
The coupling of the conservative and stochastic terms maintains a relatively high temperature $\kappa_{h}^{-1}$, and a large friction coefficient $\gamma$ is applied such that the distribution is almost always Gaussian during the evolution. As shown in the following section, the distribution of walkers $q_{t}(\mathbf{z})$ converges to $\propto\exp(-\kappa_{h}G(\mathbf{z}))$ characterized by $\mathbf{m}_{\infty}$ and $V_{\infty}$. Accordingly, the balance between exploitation and exploration is controlled using two temperatures $\kappa_{l}^{-1}$ and $\kappa_{h}^{-1}$. As $\kappa_{l}^{-1}$ decreases, the random walker distribution concentrates near the max-residual points of the current model, reflecting the role of exploitation. Conversely, as $\kappa_{h}^{-1}$ increases, the distribution smoothens progressively, enhancing the exploration of the uncharted regions.
Adaptive parameter estimation and FES construction
In this section, we show that the sampling distribution governed by Eq. (11) converges to a steady state as an approximation of $q^{\ast}$ with the proper choices of $\mathbf{m}_{t}$ and $V_{t}$ given by
$\mathbf{m}_{t}$ $=\sum_{i=1}^{N_{w}}\mathbf{z}^{i}_{t}\hat{p}(\mathbf{z}^{i}_{t}),$ (12)
$V_{t}$ $=\kappa_{t}\sum_{i=1}^{N_{w}}(\mathbf{z}^{i}_{t}-\mathbf{m}_{t})(\mathbf{z}^{i}_{t}-\mathbf{m}_{t})^{T}\hat{p}(\mathbf{z}^{i}_{t}),$
@@
where $\hat{p}(\mathbf{z})$ is defined by (10). In particular, with $\kappa_{t}=\kappa_{l}+\kappa_{h}$, (11) converges the target residue-based distribution.
Proposition 2.
@@
Suppose $\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})$ takes a local quadratic approximation in form of $\frac{1}{2}(\mathbf{z}-\boldsymbol{\mu})^{T}\Sigma^{-1}(\mathbf{z}-\boldsymbol{\mu})$, $q_{t}\to\frac{\exp{(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))}}{\int\exp{(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))}\mathrm{d}\mathbf{z}}$ as $t\to\infty$, given $\kappa_{t}=\kappa_{l}+\kappa_{h}$.
We refer to Appendix for the proof. The stochastic dynamics (11) and (12) differs from the regular Langevin dynamics; it is nonlocal and similar to the one in Ref. [47] except that a parameter $\kappa_{h}$ appears in the target distribution that modulates the exploration of the sampling dynamics. The quadratic assumption of the loss function $\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})$ is due to the fact that we are mainly interested in the regime near the max-residue point. Under the low-temperature limit, the local regime can be well characterized by the first and second moments following Laplace’s principle. While $\mathbf{m}$ identifies the extremal point, $V$ recognizes the anisotropic nature among the different CVs. In this study, for the sake of computational efficiency, we further simplify $V$ by only considering the diagonal entries denoted as $\mathbf{v}$. Moreover, we utilize a moving average to ensure stable estimation
$\mathbf{m}_{t+1}$ $=\beta_{1}\mathbf{m}_{t}+(1-\beta_{1})\sum_{i=1}^{N_{w}}\mathbf{z}^{i}_{t}\hat{p}(\mathbf{z}^{i}_{t}),$ (13)

### v4

@@ 2.3 Exploitation and exploration in the max-problem
(10) $\lim_{\kappa\to\infty}\left(-\frac{1}{\kappa}\log\left(\int\exp{(-\kappa f(\mathbf{z}))}\mathrm{d}\rho^{*}(\mathbf{z})\right)\right)=f(\mathbf{z}^{*}),$
@@
which holds true for any compactly supported probability measure $\rho^{\ast}$, where $\mathbf{z}^{\ast}\in\text{supp}(\rho^{*})$ uniquely minimizes the function $f$. This enables us to identify the max-residual point from a collection of samples $\left\{\mathbf{z}^{i}\right\}_{i=1}^{N_{w}}$ by the first-order moment under the weighted density function $p(\mathbf{z})$, i.e.,
@@
(11) $\mathbf{m}_{\kappa_{l}}=\int\mathbf{z}\frac{p(\mathbf{z})}{\int p(\mathbf{z}^{\prime})\rho(\mathrm{d}\mathbf{z}^{\prime})}\rho(\mathrm{d}\mathbf{z})\approx\sum_{i=1}^{N_{w}}\mathbf{z}^{i}\hat{p}(\mathbf{z}^{i})\quad\hat{p}(\mathbf{z})=\frac{\exp{(-\kappa_{l}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))}}{\sum_{i=1}^{N_{w}}\exp{(-\kappa_{l}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}^{i})})}$
where $\kappa_{l}^{-1}$ represents a low temperature limit, $p(\mathbf{z})=\exp(-k_{l}\mathcal{L_{N}^{-}}(\mathbf{z}))$ and $\rho=\frac{1}{N_{w}}\sum_{i=1}^{N_{w}}\delta_{\mathbf{z}^{i}}$ is the empirical measure. In order to sample the empirical measure towards $q^{*}$, we treat each sampler $\mathbf{z}^{i}$ as a random walker $\mathbf{z}^{i}_{t}$ governed by the following McKean SDE
@@
(12) $\mathrm{d}\mathbf{z}^{i}_{t}=-\frac{1}{\gamma}\nabla_{\mathbf{z}}G(\mathbf{z}^{i}_{t})\mathrm{d}t+\sqrt{\frac{2}{\kappa_{h}\gamma}}\mathrm{d}\mathbf{W}^{i}_{t},\quad i=1,\cdots,N_{w},$
[…]
@@
The coupling of the conservative and stochastic terms maintains a relatively high temperature $\kappa_{h}^{-1}$. Also, a large friction coefficient $\gamma$ is applied such that the distribution of walkers $q_{t}(\mathbf{z})$ is almost always Gaussian during the evolution. As shown in Prop. 2.4, by properly choosing the form of $V_{\kappa_{l},t}$, the distribution $q_{t}(\mathbf{z})$ converges to $\propto\exp(-\kappa_{h}G(\mathbf{z}))$ characterized by $\mathbf{m}_{\kappa_{l},\infty}$ and $V_{\kappa_{l},\infty}$. Accordingly, the balance between exploitation and exploration is controlled using two temperatures $\kappa_{l}^{-1}$ and $\kappa_{h}^{-1}$. As $\kappa_{l}^{-1}$ decreases, the distribution concentrates near the max-residual points, reflecting the role of exploitation. Conversely, as $\kappa_{h}^{-1}$ increases, the distribution smoothens progressively, enhancing the exploration of the uncharted regions.
Remark 2.3.
The present sampling dynamics (12) takes a different form from the one proposed in Ref. [7]. Specifically, the conservative potential $G(\mathbf{z})$ is constructed by both the first and second moment which enables us to conveniently characterize $\mathcal{L_{N}}(\mathbf{z})$ near the max-residual region with Laplace’s approximation, whereas it is only determined by the first moment in Ref. [7]. Also, $\kappa_{h}$ is introduced as an independent parameter to modulate the exploration of the phase space. In principle, the consensus dynamics in Ref. [7] could be also used for adaptive sampling with the proper choice of the multiplicative noise term. We proposed the modified form (12) such that the sampling parameters can be chosen with a clear physical interpretation.
2.4 Convergence analysis
In this subsection, we analyze the long-time behavior of the sampling dynamics (12). We show that the particle distribution converges exponentially fast to a steady state as an approximation of the target residual-based distribution $q^{*}(\mathbf{z})\propto\exp(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))$ under mild conditions. Specifically, we choose $\mathbf{m}_{\kappa_{l},t}$ and $V_{\kappa_{l},t}$ in the form of
(13) $\mathbf{m}_{\kappa_{l},t}=\mathcal{M}_{\kappa_{l}}(\rho_{t})\quad V_{\kappa_{l},t}=\mathcal{V}_{\kappa_{l}}(\rho_{t}),$
where $\rho_{t}=\frac{1}{N_{w}}\sum_{i=1}^{N_{w}}\delta_{\mathbf{z}^{i}_{t}}$ is the empirical measure and
(14) $\mathcal{M}_{\kappa_{l}}(\rho)$ $=\int\mathbf{z}\frac{p(\mathbf{z})}{\int p(\mathbf{z}^{\prime})\rho(\mathrm{d}\mathbf{z}^{\prime})}\rho(\mathrm{d}\mathbf{z}),$
$\mathcal{V}_{\kappa_{l}}(\rho)$ $=\kappa_{t}\int\mathbf{(}\mathbf{z}-\mathcal{M}_{\kappa_{l}}(\rho))\otimes(\mathbf{z}-\mathcal{M}_{\kappa_{l}}(\rho))\frac{p(\mathbf{z})}{\int p(\mathbf{z}^{\prime})\rho(\mathrm{d}\mathbf{z}^{\prime})}\rho(\mathrm{d}\mathbf{z}).$
@@ 2.4 Convergence analysis
In particular, we show that by choosing $\kappa_{t}=\kappa_{l}+\kappa_{h}$, Eq. (12) converges to the target distribution $q^{*}(\mathbf{z})$.
Proposition 2.4.
@@
Suppose $\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})$ takes a local quadratic approximation in form of $\frac{1}{2}(\mathbf{z}-\mathbf{\mu})^{\top}\Sigma^{-1}(\mathbf{z}-\mathbf{\mu})$. If the dynamics converge to an invariant distribution, then the stationary density is given by
(15) $q_{\infty}=\frac{\exp{(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))}}{\int\exp{(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))}\mathrm{d}\mathbf{z}},$
by choosing $\kappa_{t}=\kappa_{l}+\kappa_{h}$.
Proof 2.5.
Let $q_{\infty}(\mathbf{z})$ denote the invariant distribution of Eq. (12). Then $q_{\infty}(\mathbf{z})$ must be the invariant distribution of the following SDE
(16) $\mathrm{d}\mathbf{z}=-\frac{1}{\gamma}V^{-1}_{\kappa_{l},\infty}(\mathbf{z}-\mathbf{m}_{\kappa_{l},\infty})\mathrm{d}t+\sqrt{\frac{2}{\gamma\kappa_{h}}}\mathrm{d}\mathbf{W}_{t},$
Since $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=\frac{1}{2}(\mathbf{z}-\mathbf{\mu})^{\top}\Sigma^{-1}(\mathbf{z}-\mathbf{\mu})$ is quadratic, the re-weighted density of a Gaussian distribution $q(\mathbf{z})\sim\mathcal{N}(\mathbf{m},V)$ remains Gaussian, i.e.,
$q(\mathbf{z})e^{-\kappa_{l}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})}\propto~{}\mathcal{N}({\mathbf{m}}_{\kappa_{l}},{V}_{\kappa_{l}}),$
where ${\mathbf{m}}_{\kappa_{l}}$ and ${V}_{\kappa_{l}}$ are defined by
(17) $\begin{split}\mathbf{m}_{\kappa_{l}}(\mathbf{m},V)&=(V^{-1}+\kappa_{l}\Sigma^{-1})^{-1}(\kappa_{l}\Sigma^{-1}\mathbf{\mu}+V^{-1}\mathbf{m}),\\ V_{\kappa_{l}}(\mathbf{m},V)&=(V^{-1}+\kappa_{l}\Sigma^{-1})^{-1}.\end{split}$
Therefore, the mean and covariance of the steady-state Gaussian distribution satisfies
$\mathbf{m}_{\kappa_{l},\infty}$ $=(\kappa_{h}V_{\kappa_{l},\infty}^{-1}+\kappa_{l}\Sigma^{-1})(\kappa_{l}\Sigma^{-1}\mathbf{\mu}+\kappa_{h}V_{\kappa_{l},\infty}^{-1}\mathbf{m}_{\kappa_{l},\infty}),$
$V_{\kappa_{l},\infty}$ $=\kappa_{t}(\kappa_{h}V_{\kappa_{l},\infty}^{-1}+\kappa_{l}\Sigma^{-1})^{-1}.$
It is easy to show that by choosing $\kappa_{t}=\kappa_{l}+\kappa_{h}$, $\mathbf{m}_{\kappa_{l},\infty}$ and $V_{\kappa_{l},\infty}$ recovers $\mathbf{\mu}$ and $\Sigma$, respectively, and the invariant distribution takes the form
$q_{\infty}(\mathbf{z})\sim\mathcal{N}\left(\mathbf{\mu},\kappa_{h}^{-1}\Sigma\right).$
We emphasize that Eqs. (12) and (13) differ from the standard Langevin dynamics; the sampling relies on neither the explicit knowledge of the target distribution $q^{*}(\mathbf{z})\propto\exp(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))$ nor the numerical evaluation of $\nabla_{\mathbf{z}}q^{*}(\mathbf{z})$ for individual sampling points. The quadratic assumption of the loss function $\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})$ is due to the fact that we are mainly interested in the thermodynamically accessible region near the max-residual point. Under the low-temperature limit, the local regime can be well characterized by the first and second moment following Laplace’s principle.
Next, we show that, under appropriate conditions, the sampling dynamics (12) converges to the target distribution exponentially fast. In particular, we consider the time discretization given by
(18) $\mathbf{z}^{i}_{t+1}=\mathbf{z}^{i}_{t}-\frac{\kappa_{t}^{-1}\mathcal{V}_{\kappa_{l}}^{-1}(\rho_{t})}{\gamma}\left(\mathbf{z}^{i}_{t}-\mathcal{M}_{\kappa_{l}}(\rho_{t})\right)\delta t+\sqrt{\frac{2\delta t}{\gamma\kappa_{h}}}\eta^{i}_{t},$
where $\eta^{i}_{t}$ are independent $\mathcal{N}(0,\mathbb{I}_{M})$ random variables and $\delta t$ is the time step.
Lemma 2.6.
If the initial law $\rho_{0}$ for (18) is Gaussian, so is the law for any $t\in\mathbb{N}_{>0}$. Also, its evolution is characterized by the first and second moments $(\mathbf{m}_{t},V_{t})$ of $\rho_{t}$, which are governed by the recurrence relation
(19) $\mathbf{m}_{t+1}$ $=\left(\mathbb{I}_{M}-\frac{\delta t\kappa_{t}^{-1}V_{k_{l}}^{-1}(\mathbf{m}_{t},V_{t})}{\gamma}\right)\mathbf{m}_{t}+\frac{\delta t\kappa_{t}^{-1}V^{-1}_{k_{l}}(\mathbf{m}_{t},V_{t})}{\gamma}\mathbf{m}_{k_{l}}(\mathbf{m}_{t},V_{t}),$
$V_{t+1}$ $=\left(\mathbb{I}_{M}-\frac{\delta t\kappa_{t}^{-1}V_{k_{l}}^{-1}(\mathbf{m}_{t},V_{t})}{\gamma}\right)^{2}V_{t}+\frac{2\delta t}{\gamma\kappa_{h}},$
where $\mathbf{m}_{k_{l}}(\mathbf{m},V)=\mathcal{M}_{\kappa_{l}}(g(\cdot;\mathbf{m},V))$ and $V_{k_{l}}(\mathbf{m},V)=\mathcal{V}_{\kappa_{l}}(g(\cdot;\mathbf{m},V))$.
Proof 2.7.
The time evolution of the law of the solution (18) is governed by the following discrete-time dynamics on probability density:
(20) $\rho_{t+1}(\mathbf{z})=\int_{\mathbb{R}^{M}}g\left(\mathbf{z};\mathbf{z}^{\prime}-\frac{\delta t\mathcal{\kappa}_{t}^{-1}V_{\kappa_{l}}^{-1}(\rho_{t})}{\gamma}\left(\mathbf{z}^{\prime}-\mathcal{M}_{\kappa_{l}}(\rho_{t})\right),\frac{2\delta t}{\gamma\kappa_{h}}\right)\rho_{t}(\mathbf{z}^{\prime})\mathrm{d}\mathbf{z}^{\prime},$
where $g(\mathbf{z};\mathbf{m},V)=\frac{1}{\sqrt{(2\pi)^{M}\det(V)}}\exp(-\frac{1}{2}\left(\mathbf{z}-\mathbf{m})^{\top}V^{-1}(\mathbf{z}-\mathbf{m})\right)$. Since both $\rho_{t}$ and $g$ are Gaussian, the convolution in (20) results in another Gaussian distribution. Therefore, $\rho_{t}$ remains Gaussian for all $t>0$. The update rules for the mean and covariance can be derived by explicitly computing the first and second moments of $\rho_{t+1}$ from the integral form (20).
Proposition 2.8.
Assume that $\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})$ takes a local quadratic approximation in form of $\frac{1}{2}(\mathbf{z}-\mathbf{\mu})^{\top}\Sigma^{-1}(\mathbf{z}-\mathbf{\mu})$ and initial condition $(\mathbf{m}_{0},V_{0})\in\mathbb{R}^{M}\times S^{M}_{++}$, where $S^{M}_{++}$ denotes the set of symmetric strictly positive definite matrices in $\mathbb{R}^{M\times M}$. Then the first and second moment of the empirical distribution $\mathbf{m}_{t}$ and $V_{t}$ converge to $\mu$ and $\kappa_{h}^{-1}\Sigma$ exponentially fast if $\delta t$ is sufficiently small. To be more specific, we have
(21) $\mathbf{m}_{t}$ $=\mu+\left(\mathbb{I}_{M}-\frac{\delta t\kappa_{t}^{-1}\kappa_{l}\Sigma^{-1}}{\gamma}\right)^{t}\mathbf{(}\mathbf{m}_{0}-\mu)$
$V_{t}$ $=\kappa_{h}^{-1}\Sigma+\Sigma\left(\mathbb{I}_{M}-\frac{2\delta t\kappa_{l}\Sigma^{-1}}{\gamma(\kappa_{l}+\kappa_{h})}\right)^{t}\left(\Sigma^{-1}V_{0}-\kappa_{h}^{-1}\mathbb{I}_{M}\right).$
Proof 2.9.
For the first moment, we have the following recurrence relation
(22) $\mathbf{m}_{t+1}$ $=\mathbf{m}_{t}-\frac{\delta t\kappa_{t}^{-1}(V^{-1}_{t}+\kappa_{l}\Sigma^{-1})}{\gamma}\left(\mathbf{m}_{t}-(V_{t}^{-1}+\kappa_{l}\Sigma^{-1})^{-1}(\kappa_{l}\Sigma^{-1}\mathbf{\mu}+V^{-1}_{t}\mathbf{m}_{t})\right)$
$=\left(\mathbb{I}_{M}-\frac{\delta t\kappa_{t}^{-1}\kappa_{l}\Sigma^{-1}}{\gamma}\right)\mathbf{m}_{t}+\frac{\delta t\kappa_{t}^{-1}\kappa_{l}\Sigma^{-1}}{\gamma}\mu.$
Therefore, we have that
(23) $\mathbf{m}_{t+1}-\mu$ $=\left(\mathbb{I}_{M}-\frac{\delta t\kappa_{t}^{-1}\kappa_{l}\Sigma^{-1}}{\gamma}\right)\mathbf{(}\mathbf{m}_{t}-\mu)$
$=\left(\mathbb{I}_{M}-\frac{\delta t\kappa_{t}^{-1}\kappa_{l}\Sigma^{-1}}{\gamma}\right)^{t+1}\mathbf{(}\mathbf{m}_{0}-\mu).$
By chossing the timestep $\delta t\in\left(0,\frac{\gamma}{\kappa_{t}^{-1}\kappa_{l}e(\Sigma)^{-1}}\right)$, where $e(\Sigma)$ denote the smallest eigenvalue of $\Sigma$, the first moment will converge to $\mu$ exponentially fast. For the second moment, we have that
(24) $V_{t+1}=\left(\mathbb{I}_{M}-\frac{\delta t\kappa_{t}^{-1}(V^{-1}_{t}+\kappa_{l}\Sigma^{-1})}{\gamma}\right)^{2}V_{t}+\frac{2\delta t}{\gamma\kappa_{h}}$
By assuming that $\delta t$ is small, we have the following approximation
(25) $V_{t+1}$ $=\left(\mathbb{I}_{M}-\frac{2\delta t\kappa_{t}^{-1}(V^{-1}_{t}+\kappa_{l}\Sigma^{-1})}{\gamma}\right)V_{t}+\frac{2\delta t}{\gamma\kappa_{h}}$
$=V_{t}-\frac{2\delta t\kappa_{l}\Sigma^{-1}V_{t}}{\gamma(\kappa_{l}+\kappa_{h})}+\frac{2\delta t\kappa_{l}}{\gamma\kappa_{h}(\kappa_{l}+\kappa_{h})}$
$=V_{t}-\frac{2\delta t\kappa_{l}}{\gamma(\kappa_{l}+\kappa_{h})}\left(\Sigma^{-1}V_{t}-\kappa_{h}^{-1}\mathbb{I}_{M}\right).$
Therefore, we have that
(26) $\Sigma^{-1}V_{t+1}-\kappa_{h}^{-1}\mathbb{I}_{M}$ $=\left(\mathbb{I}_{M}-\frac{2\delta t\kappa_{l}\Sigma^{-1}}{\gamma(\kappa_{l}+\kappa_{h})}\right)\left(\Sigma^{-1}V_{t}-\kappa_{h}^{-1}\mathbb{I}_{M}\right)$
$=\left(\mathbb{I}_{M}-\frac{2\delta t\kappa_{l}\Sigma^{-1}}{\gamma(\kappa_{l}+\kappa_{h})}\right)^{t+1}\left(\Sigma^{-1}V_{0}-\kappa_{h}^{-1}\mathbb{I}_{M}\right).$
By choosing the timestep $\delta t\in\left(0,\frac{\gamma}{2\kappa_{t}^{-1}\kappa_{l}e(\Sigma)^{-1}}\right)$, the second moment will converge to $\kappa_{h}^{-1}\Sigma$ exponentially fast.
2.5 Practical stabilization modifications
In this study, for the sake of computational efficiency, we further simplify $V$ by only considering the diagonal entries denoted as $\mathbf{v}$. Moreover, we utilize a moving average to ensure stable estimation
(27) $\mathbf{m}_{t+1}$ $=\beta_{1}\mathbf{m}_{t}+(1-\beta_{1})\sum_{i=1}^{N_{w}}\mathbf{z}^{i}_{t}\hat{p}(\mathbf{z}^{i}_{t}),$